OpenStudy (sissyedgar):

uhh... hi! :P ok luk in comments 4 file! :)

sammixboo (sammixboo):

Weee

OpenStudy (sissyedgar):

:P

OpenStudy (alexandervonhumboldt2):

q=-12*4/3

OpenStudy (alexandervonhumboldt2):

what is -12*4 and then divided by 3?

OpenStudy (sissyedgar):

-16?

sammixboo (sammixboo):

-3/4 = -0.75, so -0.75q = 12 you solve this by dividing -0.75 on both sides of the equation -0.75q/-0.75 = 12/-0.75

sammixboo (sammixboo):

Correct =3

OpenStudy (alexandervonhumboldt2):

-48/3=-16 correct!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (sissyedgar):

:D